(19) Show that if n is an integer then n^2 ≡ 0 or 1 (mod 4).
(20) Use the result of Exercise (19) to show that if m is a positive integer that can be written in the form m= 4k+ 3 (where k is a nonnegative integer), then m is not the sum of the squares of two integers.
1
Expert's answer
2020-10-26T19:21:13-0400

(19) Ans : Any integer "n" under( mod 4 ) is

"n\equiv"0 (mod 4) or "n\equiv" 1 (mod 4 ) or "n\equiv" 2 (mod 4) or "n\equiv" 3 (mod 4)

Again we know that from property of congruences that

If "a\equiv b" (mod n) , then "a^k\equiv b^k" (mod n) for any positive integer "k" .

Therefore , "n^2\equiv 0^2" (mod 4) "\implies" "n^2\equiv 0" (mod 4),

"n^2\equiv 1^2" (mod 4) "\implies n^2\equiv 1" (mod 4),

"n^2\equiv2^2" (mod 4) "\implies n^2 \equiv 4" (mod 4) "\implies n^2 \equiv 1" (mod 4),

"n^2\equiv 3^2" (mod 4) "\implies n^2 \equiv 9" (mod 4) "\implies n^2 \equiv 1" (mod 4)

Hence , for any integer n , "n^2\equiv 0 \ or\ 1" (mod 4)

(20) Ans : Given that "m" is a positive integer in the form "4k+3" , i,e "m=4k+3" for some integer "k \in \Z_{\geq 0}".

Claim : "m" can not be written as a sum of square of two integer .

If possible let "m=a^2+b^2" , where "a" and "b" are two integer .

We know that from (19) , for any integer "n" , "n^2\equiv 0 \ or \ 1" (mod 4)

Therefore , the possible value of "(a^2+b^2)" (mod 4 ) are 0 ,1 and 2 .

But "m" (mod 4 )= 3 , which is a contradiction of "m=a^2+b^2" .

Hence , m is not the sum of square of two integer .